A Multiple-choice test has 15 questions with each having 4 possible answers with one correct. Assume a student randomly guesses the answer to every question.
a. What is the probability of getting exactly 5 correct answers?

b. What is the probability of getting less than 6 correct answers?
I have to have this in by TODAY - August 6th, 2010. Please help!

Answer by stanbon(75887)

A Multiple-choice test has 15 questions with each having 4 possible answers with one correct. Assume a student randomly guesses the answer to every question.
---
Binomial Problem with n=15 and p = 1/4
--------
a. What is the probability of getting exactly 5 correct answers?
P(x= 5) = 15C5*(1/4)^5*(3/4)^10 = binompdf(15,1/4,5) = 0.1651
-------------------------
b. What is the probability of getting less than 6 correct answers?
P(less than 6) = binomcdf(15,1/4,5) = 0.8516
=======================
Cheers,
Stan H.